# Adversarial Defence without Adversarial Defence: Enhancing Language Model Robustness via Instance-level Principal Component Removal

### v1

@@ 1 Introduction
Existing adversarial defence methods often demand extensive computational resources, or have limited improvements in adversarial robustness. For example, adversarial training-based methods Madry et al. (2018); Zhu et al. (2020); Li and Qiu (2021); Wang et al. (2021a) involve generating perturbations through multiple iterations during training, which significantly increases the computational overhead. Similarly, some ensemble-based techniques leverage the statistical properties of the ensemble to provably certify the robustness Ye et al. (2020a); Zhou et al. (2021b); Moon et al. (2023b); Zeng et al. (2023), leading to additional costs during both training and inference. An alternative line of defence leverages regularisation-based methods Ishida et al. (2020); Wang et al. (2021a); Liu et al. (2022); Yang et al. (2023b), which are more computationally efficient but tend to show limited improvements in robustness against adversarial attacks Zhu and Rao (2023). This disparity highlights the need for more efficient adversarial defence approaches that strike a balance between computational efficiency and robustness enhancement.
To address these challenges, we propose Purified Representation (PuRe) to enhance adversarial robustness without introducing adversarial perturbations during training, either implicitly or explicitly [^1] The title “Adversarial Defence without Adversarial Defence” is intended to be paradoxical. It signifies that PuRe enhances adversarial robustness without employing the common strategies of adversarial defence (e.g., generating adversarial examples for training).. PuRe is implemented as a module that is integrated directly into the PLM’s architecture. The entire model is then trained using a standard fine-tuning process, requiring no special modifications. At its core, this module leverages Principal Component Removal Arora et al. (2017) to reshape the embedding space. By removing dominant components, it encourages representations to align more closely with Gaussian-like distributions, which reduces the model’s sensitivity to the targeted perturbations that adversaries often exploit. This transformation strengthens robustness without relying on adversarial example generation or resource-intensive training augmentations, providing an efficient and practical solution for improving adversarial resilience in NLP tasks. The evaluation of PuRe is underpinned by benchmarking eight language understanding datasets, spanning across sentiment analysis, subjectivity status classification, paraphrase identification, textual entailment, and commonsense reasoning. PuRe shows superior textual adversarial defence ability to most tasks, while performing on-par with the baselines in terms of before-attack accuracy, indicating a good trade-off between robustness and generalisation. Our contributions can be summarised as follows:
@@ 2.1 Adversarial Attacks
Transferring adversarial attack methods from computer vision to NLP introduces unique challenges due to the discrete nature of textual data as opposed to continuous pixel values. Thus, NLP-focused adversarial attack research has largely focused on crafting semantics-preserving perturbations. For example, back-translation Iyyer et al. (2018) generates adversarial examples by translating text back and forth between different languages. Wang et al. (2020b) use GANs Goodfellow et al. (2014a) to create fluent adversarial texts that closely resemble natural language. Additionally, methods have been developed to identify critical words in text and replace them with synonyms or to introduce character-level perturbations such as typos in letters, numbers, or special symbols Jin et al. (2020); Maheshwary et al. (2021); Li et al. (2018).
@@ 2.2 Adversarial Defences
Adversarial training-based methods  involve augmenting the training data with adversarial examples, enabling the model to learn in an environment that simulates attacks in the training process, either implicitly or explicitly Jin et al. (2020); Morris et al. (2020); Si et al. (2020); Wang et al. (2021a); Hauser et al. (2023). Implicit approaches usually generate perturbations dynamically in the embedding space as a part of the training process, which improves the model’s resilience to a range of adversarial scenarios Wu et al. (2017); Zhu et al. (2020); Dong et al. (2021a); Gao et al. (2023); Latorre et al. (2023). Explicit approaches, on the other hand, involves generating adversarial examples in the input space (text data) using adversarial attack methods Jin et al. (2020); Li et al. (2020c); Tan et al. (2020); Zang et al. (2020), and these pre-generated adversarial examples will be incorporated into the training pipeline. We refer this explicit adversarial training-based approach as Adversarially-augmented (AdvAug) training (see §5.2). Despite its efficacy and interpretability, the adversarial training-based methods are often computationally intensive due to the need for extensive adversarial example generation and fine-tuning.
Perturbation control-based methods  aim to detect and correct adversarial inputs by incorporating mechanisms to recognise potential perturbations Alshemali and Kalita (2019); Yoo et al. (2022); Shen et al. (2023); Ali et al. (2023) or by altering the perturbation toward cleaner inputs to limit the adversarial space Sato et al. (2018); Zhang et al. (2020); Zhou et al. (2021a); Bao et al. (2021). Techniques include spell-checking systems for character-level defences that correct adversarially manipulated inputs before classification Alshemali and Kalita (2019) and word-level defences that substitute input words with synonyms to neutralise adversarial effects Ye et al. (2020a); Zhou et al. (2021a); Dong et al. (2021b). However, synonym-based methods often face limitations in practical scenarios, where the perturbation sets of potential attacks are usually unknown Li et al. (2021a).
Certification-based methods  provide theoretical guarantees by constructing a perturbation-resistant region around the input space Wang et al. (2019); Dong et al. (2021b); Asl et al. (2023); Moon et al. (2023a); Zeng et al. (2023). Although these methods offer strong theoretical assurances, they typically involve impractical constraints in real-world applications. Certification-based methods can require extensive computational resources and long verification times Zeng et al. (2023), which may not be feasible in applications with limited computational capacity or real-time processing requirements.
@@ 4.4 Datasets
We evaluate PuRe across eight language understanding datasets covering various NLP tasks such as: sentiment analysis, subjectivity status classification, paraphrase identification, textual entailment, and commonsense reasoning. In contrast to other studies Dong et al. (2021b); Bao et al. (2021); Li et al. (2021a); Wang et al. (2022a); Shen et al. (2023); Hu et al. (2023); Zeng et al. (2023); Zhan et al. (2023); Moon et al. (2023b), which often restrict their evaluations to a limited selection of test samples from their datasets, we extend our analysis to include the entire test sets for all eight datasets, ensuring a comprehensive assessment. This broad evaluation approach contrasts with the common practice in the field, where researchers only utilise a small portion of available test data, which may not fully represent the model’s performance across different scenarios.
@@ 4.5 Implementation Details
[…]
@@ 5.8 Run Time Analysis
Fine-tune 1.01.01.0 1.01.01.0 -
PGD Madry et al. (2019) ×3.3$\times$ 3.3× 3.3 ×1.0$\times$ 1.0× 1.0 10.8210.8210.82
FreeLB Zhu et al. (2020) ×2.6$\times$ 2.6× 2.6 ×1.0$\times$ 1.0× 1.0 12.0712.0712.07
InfoBERT Wang et al. (2020a) ×3.8$\times$ 3.8× 3.8 ×1.0$\times$ 1.0× 1.0 4.544.544.54
TAVAT Li and Qiu (2020) ×1.6$\times$ 1.6× 1.6 ×1.0$\times$ 1.0× 1.0 10.4310.4310.43
SAFER Ye et al. (2020b) ×1.1$\times$ 1.1× 1.1 ×1.0$\times$ 1.0× 1.0 16.9916.9916.99
DNE Zhou et al. (2021b) ×2.5$\times$ 2.5× 2.5 ×3.0$\times$ 3.0× 3.0 0.450.450.45
Flooding-X Liu et al. (2022) ×1.0$\times$ 1.0× 1.0 ×1.0$\times$ 1.0× 1.0 3.813.813.81
ALS Yang et al. (2023b) ×1.1$\times$ 1.1× 1.1 ×1.0$\times$ 1.0× 1.0 3.703.703.70
AdvFooler Hoang et al. (2024) ×1.0$\times$ 1.0× 1.0 ×1.5$\times$ 1.5× 1.5 3.713.713.71
PuRe (w/ SVD) ×1.9$\times$ 1.9× 1.9 ×1.5$\times$ 1.5× 1.5 13.6613.6613.66
PuRe (w/ rSVD) ×1.2$\times$ 1.2× 1.2 ×1.1$\times$ 1.1× 1.1 17.1717.1717.17
Table 7: Runtime comparison of PuRe and baseline methods on the MRPC dataset, with $\Delta$Apdr indicating the absolute drop in Apdr relative to the fine-tuning baseline.
@@ Acknowledgment
We are grateful to the action editor and the anonymous reviewers for their thoughtful feedback and valuable suggestions, which have helped improve the quality and clarity of this work. This work was partially supported by the Economic and Social Research Council (ES/V011278/1) and Engineering and Physical Sciences Research Council (EP/Y009800/1).
@@ Appendix C Adversarial Defence Baselines
[…]

### v2

@@ 1 Introduction
Existing adversarial defence methods often demand extensive computational resources, or have limited improvements in adversarial robustness. For example, adversarial training-based methods Madry et al. (2018); Zhu et al. (2020); Li and Qiu (2021); Wang et al. (2021a) involve generating perturbations through multiple iterations during training, which significantly increases the computational overhead. Similarly, some ensemble-based techniques leverage the statistical properties of the ensemble to provably certify the robustness Ye et al. (2020a); Zhou et al. (2021b); Moon et al. (2023b); Zeng et al. (2023), leading to additional costs during both training and inference. An alternative line of defence leverages regularisation-based methods Ishida et al. (2020); Wang et al. (2021a); Liu et al. (2022); Yang et al. (2023b), which are more computationally efficient but tend to show limited improvements in robustness against adversarial attacks Zhu and Rao (2023). This disparity highlights the need for more efficient defence methods that strike a balance between computational efficiency and robustness enhancement.
To address these challenges, we propose Purified Representation (PuRe) to enhance adversarial robustness without introducing adversarial perturbations during training, either implicitly or explicitly [^1] The title is intended to be paradoxical. It signifies that PuRe enhances adversarial robustness without employing the common strategies of adversarial defence (e.g., generating adversarial examples for training).. PuRe is implemented as a module that is integrated directly into the PLM’s architecture. The entire model is then trained using a standard fine-tuning process, requiring no special modifications. At its core, this module leverages Principal Component Removal Arora et al. (2017) to reshape the embedding space. By removing dominant components, it encourages representations to align more closely with Gaussian-like distributions, which reduces the model’s sensitivity to the targeted perturbations that adversaries often exploit. This transformation strengthens robustness without relying on adversarial example generation or resource-intensive training augmentations, providing an efficient and practical solution for improving adversarial resilience in NLP tasks. The evaluation of PuRe is underpinned by benchmarking eight language understanding datasets, spanning across sentiment analysis, subjectivity status classification, paraphrase identification, textual entailment, and commonsense reasoning. PuRe shows superior textual adversarial defence ability to most tasks, while performing on-par with the baselines in terms of before-attack accuracy, indicating a good trade-off between robustness and generalisation. Our contributions can be summarised as follows:
@@ 2.1 Adversarial Attacks
Transferring adversarial attack methods from computer vision to NLP introduces unique challenges due to the discrete nature of textual data as opposed to continuous pixel values. Thus, NLP-focused adversarial attack research has largely focused on crafting semantics-preserving perturbations. For example, back-translation Iyyer et al. (2018) generates adversarial examples by translating text back and forth between different languages. Wang et al. (2020b) use GANs Goodfellow et al. (2014a); Li et al. (2020d) to create fluent adversarial texts that closely resemble natural language. Additionally, methods have been developed to identify critical words in text and replace them with synonyms or to introduce character-level perturbations such as typos in letters, numbers, or special symbols Jin et al. (2020); Maheshwary et al. (2021); Li et al. (2018).
@@ 2.2 Adversarial Defences
Adversarial training-based methods  involve augmenting the training data with adversarial examples, enabling the model to learn in an environment that simulates attacks in the training process, either implicitly or explicitly Jin et al. (2020); Morris et al. (2020); Si et al. (2020); Wang et al. (2021a); Hauser et al. (2023). Implicit approaches usually generate perturbations dynamically in the embedding space as a part of the training process, which improves the model’s resilience to a range of adversarial scenarios Wu et al. (2017); Zhu et al. (2020); Dong et al. (2021); Gao et al. (2023); Latorre et al. (2023). Explicit approaches, on the other hand, involves generating adversarial examples in the input space (text data) using adversarial attack methods Jin et al. (2020); Li et al. (2020c); Tan et al. (2020); Zang et al. (2020), and these pre-generated adversarial examples will be incorporated into the training pipeline. We refer this explicit adversarial training-based approach as Adversarially-augmented (AdvAug) training (see §5.2). Despite its efficacy and interpretability, the adversarial training-based methods are often computationally intensive due to the need for extensive adversarial example generation and fine-tuning.
Perturbation control-based methods  aim to detect and correct adversarial inputs by incorporating mechanisms to recognise potential perturbations Alshemali and Kalita (2019); Yoo et al. (2022); Shen et al. (2023); Ali et al. (2023) or by altering the perturbation toward cleaner inputs to limit the adversarial space Sato et al. (2018); Zhang et al. (2020); Zhou et al. (2021a); Bao et al. (2021). Techniques include spell-checking systems for character-level defences that correct adversarially manipulated inputs before classification Alshemali and Kalita (2019) and word-level defences that substitute input words with synonyms to neutralise adversarial effects Ye et al. (2020a); Zhou et al. (2021a); Dong et al. (2021). However, synonym-based methods often face limitations in practical scenarios, where the perturbation sets of potential attacks are usually unknown Li et al. (2021a).
Certification-based methods  provide theoretical guarantees by constructing a perturbation-resistant region around the input space Wang et al. (2019); Dong et al. (2021); Asl et al. (2023); Moon et al. (2023a); Zeng et al. (2023). Although these methods offer strong theoretical assurances, they typically involve impractical constraints in real-world applications. Certification-based methods can require extensive computational resources and long verification times Zeng et al. (2023), which may not be feasible in applications with limited computational capacity or real-time processing requirements.
@@ 4.4 Datasets
We evaluate PuRe across eight language understanding datasets covering various NLP tasks such as: sentiment analysis, subjectivity status classification, paraphrase identification, textual entailment, and commonsense reasoning. In contrast to other studies Dong et al. (2021); Bao et al. (2021); Li et al. (2021a); Wang et al. (2022a); Shen et al. (2023); Hu et al. (2023); Zeng et al. (2023); Zhan et al. (2023); Moon et al. (2023b), which often restrict their evaluations to a limited selection of test samples from their datasets, we extend our analysis to include the entire test sets for all eight datasets, ensuring a comprehensive assessment. This broad evaluation approach contrasts with the common practice in the field, where researchers only utilise a small portion of available test data, which may not fully represent the model’s performance across different scenarios.
@@ 4.5 Implementation Details
To keep experiments simple and reproducible, we avoid extensive hyper-parameter tuning and instead apply a light grid search over a small set of commonly used values: batch sizes $\{8,16,32\}$ and learning rates $\{1\mathrm{e}{-5},2\mathrm{e}{-5},5\mathrm{e}{-5}\}$. For the three adversarial attackers, all the four evaluation metrics are tested on the entire test set for every dataset on sequence classification tasks.
@@ 5.8 Run Time Analysis
Fine-tune $1.0$ $1.0$ -
PGD Madry et al. (2019) $\times 3.3$ $\times 1.0$ $10.82$
FreeLB Zhu et al. (2020) $\times 2.6$ $\times 1.0$ $12.07$
InfoBERT Wang et al. (2020a) $\times 3.8$ $\times 1.0$ $4.54$
TAVAT Li and Qiu (2020) $\times 1.6$ $\times 1.0$ $10.43$
SAFER Ye et al. (2020b) $\times 1.1$ $\times 1.0$ $16.99$
DNE Zhou et al. (2021b) $\times 2.5$ $\times 3.0$ $0.45$
Flooding-X Liu et al. (2022) $\times 1.0$ $\times 1.0$ $3.81$
ALS Yang et al. (2023b) $\times 1.1$ $\times 1.0$ $3.70$
AdvFooler Hoang et al. (2024) $\times 1.0$ $\times 1.5$ $3.71$
PuRe (w/ SVD) $\times 1.9$ $\times 1.5$ $13.66$
PuRe (w/ rSVD) $\times 1.2$ $\times 1.1$ $17.17$
[…]
@@ Acknowledgment
We are grateful to the action editor and the anonymous reviewers for their thoughtful feedback and valuable suggestions, which have helped improve the quality and clarity of this work. Stuart E. Middleton was supported by the Economic and Social Research Council (ES/V011278/1) and Engineering and Physical Sciences Research Council (EP/Y009800/1).
@@ Appendix C Adversarial Defence Baselines
where $\nabla_{\delta}\mathcal{L}\left(f_{\theta}(x+\delta^{(t)}),y\right)$ is the gradient of the loss with respect to $\delta$, and $\text{sgn}(\cdot)$ takes the sign of each component in the gradient.